# A Survey on Mixture of Experts

### v1

@@ Abstract.
Large language models (LLMs) have garnered unprecedented advancements across diverse fields, ranging from natural language processing to computer vision and beyond. The prowess of LLMs is underpinned by their substantial model size, extensive and diverse datasets, and the vast computational power harnessed during training, all of which contribute to the emergent abilities of LLMs (e.g., in-context learning) that are not present in small models. Within this context, the mixture of experts (MoE) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead, gaining significant attention from academia and industry. Despite its growing prevalence, there lacks a systematic and comprehensive review of the literature on MoE. This survey seeks to bridge that gap, serving as an essential resource for researchers delving into the intricacies of MoE. We first briefly introduce the structure of the MoE layer, followed by proposing a new taxonomy of MoE. Next, we overview the core designs for various MoE models including both algorithmic and systemic aspects, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, we delineate the multifaceted applications of MoE in practice, and outline some potential directions for future research. To facilitate ongoing updates and the sharing of cutting-edge developments in MoE research, we have established a resource repository accessible at https://github.com/withinmiaov/A-Survey-on-Mixture-of-Experts.
1. Introduction
@@
In the current landscape of artificial general intelligence (AGI), the transformative impact of transformer-based large language models (LLMs) has permeated diverse fields such as natural language processing (Vaswani et al., 2017; Brown et al., 2020; Wei et al., 2022b; Ouyang et al., 2022; Chowdhery et al., 2023; Achiam et al., 2023; Jiang et al., 2024c), computer vision (Riquelme et al., 2021; Liu et al., 2021), and multimodality (Lu et al., 2019; Zhang et al., 2021; Zhou et al., 2022b; Zhu et al., 2023a). Building upon the foundational transformer architecture, LLMs demonstrate extraordinary capabilities, which are attributed to their sheer size, the breadth of data they are trained on, and the significant computational resources invested in their development (Kaplan et al., 2020; Wei et al., 2022a; Yoo et al., 2024). Recognizing a scaling law (Kaplan et al., 2020; Hoffmann et al., 2022) that underpins their evolution, it is imperative to identify and implement efficient methodologies for the sustainable scaling of LLMs.
@@
The concept of mixture of experts (MoE), initially introduced in (Jacobs et al., 1991; Jordan and Jacobs, 1994), has undergone extensive exploration and advancement as evidenced by subsequent studies (Collobert et al., 2001; Rasmussen and Ghahramani, 2001; Shahbaba and Neal, 2009; Eigen et al., 2013; Theis and Bethge, 2015; Deisenroth and Ng, 2015; Aljundi et al., 2017). The emergence of sparsely-gated MoE (Shazeer et al., 2017), particularly within the integration of transformer-based large language models (Lepikhin et al., 2020), has brought new vitality to this three-decade-old technology. The MoE framework is based on a simple yet powerful idea: different parts of a model, known as experts, specialize in different tasks or aspects of the data. With this paradigm, only pertinent experts are engaged for a given input, keeping the computational cost in check while still benefiting from a large pool of specialized knowledge. This scalable and flexible innovation has offered an effective approach for adhering to the scaling law, allowing for increased model capacity without a corresponding surge in computational demands. As depicted in Figure 1, MoE has maintained a robust trajectory of growth, particularly notable in 2024 with the advent of Mixtral-8x7B (Jiang et al., 2024a) and a variety of subsequent industrial-scale LLMs such as Grok-1 (xAI, 2024), DBRX (Databricks, 2024), Arctic (Team, 2024c), DeepSeek-V2 (DeepSeek-AI, 2024), etc.
@@
Despite the increasing popularity and application of MoE models in various domains, the literature has yet to see a survey that thoroughly examines and categorizes the advancements in this area. The most recent review of MoE we could find was presented in September 2022 (Fedus et al., 2022a), predating the pivotal “ChatGPT moment”, which omits the significant advancements that have recently emerged alongside the escalating academic and industrial interest in this domain. This gap in the literature not only hinders the progress of MoE research but also limits the dissemination of knowledge on this topic to a broader audience. Our survey aims to address this deficit by providing a clear and comprehensive overview of MoE with a novel taxonomy that segments recent progress into algorithm, system and application.
@@
Under this taxonomy, we first delve into MoE algorithmic advancements, particularly the prevalent substitution of feed-forward network (FFN) layers with MoE layers in transformer-based LLMs (Lepikhin et al., 2020; Du et al., 2022; Fedus et al., 2022b; Zoph et al., 2022; Jiang et al., 2024a; Xue et al., 2024; DeepSeek-AI, 2024). As each MoE layer integrates multiple FFNs—each designated as an expert—and employs a gating function to activate a select subset of these experts, we explore the design choices of gating function and expert network, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, to underscore the flexibility and versatility of MoE, we extend our analysis beyond the standard integration of MoE into model backbone, and discuss an array of novel MoE-related designs, such as soft MoE with token or expert merging (Puigcerver et al., 2023; Muqeeth et al., 2023; Zhong et al., 2024; Zadouri et al., 2023; Wu et al., 2023a), mixture of parameter-efficient experts (MoPEs) (Wang et al., 2022; Zadouri et al., 2023; Dou et al., 2023b; Gou et al., 2023; Luo et al., 2024; Wu et al., 2024a), training and inference schemes with model transition between dense and sparse (Komatsuzaki et al., 2022; Zhang et al., 2022a; Team, 2023; Xue et al., 2022a; Chen et al., 2022b; Sukhbaatar et al., 2024), and various derivatives (Chen et al., 2023c; Antoniak et al., 2023; Raposo et al., 2024; Xue et al., 2022b; Tan et al., 2023; Choi et al., 2023).
@@
With the gradual convergence of model architecture design in industrial products, system design has emerged as a pivotal factor in enhancing the quality of LLM services. Given the close association of MoE models with machine learning system design, we provide a comprehensive overview of MoE system design, including computation, communication and storage enhancements tailored to address the unique challenges posed by the sparse and dynamic nature of its computational workload. Additionally, we overview the applications of MoE across various domains, including natural language processing, computer vision, recommender system, and multimodal contexts.
@@
The remainder of this survey is organized as follows. Section 2 provides a foundational understanding of MoE, contrasting sparse and dense activation of experts. Section 3 introduces our proposed taxonomy for categorizing MoE advancements. Sections 4, 5, and 6 delve into the algorithmic designs, computing system support, and various applications of MoE models, following the structure outlined in our taxonomy in Figure 3. Finally, in Section 7, we highlight the critical challenges and opportunities for bridging the research-practicality gap, culminating in Section 8 with our conclusions.
2. Background on Mixture of Experts
@@
In transformer-based large language models (LLMs), each mixture of experts (MoE) layer typically consists of a set of “expert networks” , alongside a “gating network” . The role of the gating network, which often takes the form of a linear network with a softmax activation function, is to direct the input to the appropriate expert networks (Shazeer et al., 2017; Fedus et al., 2022b). The MoE layer is strategically placed to select the feed-forward network (FFN) within each Transformer block, typically following the self-attention (SA) sub-layer. This positioning is crucial because the FFN become increasingly computationally demanding as the model scales up. For instance, in the PaLM (Chowdhery et al., 2023) model with the parameter number of 540B, the 90% of these parameters are within its FFN layers.
@@
Formally, each expert network , usually a linear-ReLU-linear network, is parameterized by , accepting the same input and generating an output . In parallel, the gating network , parameterized by and typically consisting of a linear-ReLU-linear-softmax network, yields the output . Based on the design of gating function, MoE layers can be broadly classified into two categories as follows.
2.1. Dense MoE
@@
The dense mixture of experts layer activates all expert networks during each iteration. This strategy has been extensively employed in a range of early works, including those by (Jacobs et al., 1991; Jordan and Jacobs, 1994; Collobert et al., 2001; Rasmussen and Ghahramani, 2001; Shahbaba and Neal, 2009; Eigen et al., 2013; Theis and Bethge, 2015; Deisenroth and Ng, 2015; Aljundi et al., 2017; Ma et al., 2018). Most recent, the dense MoE concept has been revisited by studies such as EvoMoE(Nie et al., 2021), MoLE (Wu et al., 2023b), LoRAMoE (Dou et al., 2023b), and DS-MoE(Pan et al., 2024). The structure of the dense MoE layer is depicted in Figure 2(a). Consequently, the output of the dense MoE layer can be formulated as
@@ 2.2. Sparse MoE
While a dense mixture of experts typically yields higher prediction accuracy (Riquelme et al., 2021), it also incurs a significant increase in computational overhead. To address this, Shazeer et al. (Shazeer et al., 2017) introduced the sparsely-gated MoE layer, which is designed to activate only a selected subset of experts during each forward pass. This strategy achieves sparsity by computing a weighted sum of the outputs from only the top- experts, rather than aggregating the outputs from all experts. The structure of the sparse MoE layer is illustrated in Figure 2(b). Building on the framework established by (Shazeer et al., 2017), Equation (2.2) can be modified to reflect the sparsely-gated mechanism as follows:
@@
To explain, function retains only the top- entries of a vector at their original values, while setting all other entries to . Following the operation, those entries assigned become approximately zero. The hyper-parameter is selected based on the specific application, with common choices being (Fedus et al., 2022b; Clark et al., 2022) or (Lepikhin et al., 2020; Du et al., 2022; Rajbhandari et al., 2022; Zoph et al., 2022; Jiang et al., 2024a; Tianwen Wei, 2024). The addition of a noise term is a prevalent strategy for the training of a sparsely-gated MoE layer, fostering exploration among experts and enhancing the stability of MoE training (Shazeer et al., 2017; Fedus et al., 2022b).
@@
Although the sparse gate substantially expands the model’s parameter space without a corresponding increase in computational cost, it can lead to a load balancing issue. Such an issue refers to the uneven distribution of workload across experts, with some being frequently utilized and others seldom or never engaged. To address this, each MoE layer incorporates an auxiliary loss function that promotes an even distribution of tokens across experts within each batch, as described in many studies (Lepikhin et al., 2020; Fedus et al., 2022b; Du et al., 2022; Jiang et al., 2024a; Lieber et al., 2024; Dai et al., 2024; Tianwen Wei, 2024). To formulate this concept, consider a batch of queries , comprising tokens, and experts indexed from to . Following (Lepikhin et al., 2020; Fedus et al., 2022b), the auxiliary load balancing loss for the batch is defined as
@@
where represents the proportion of tokens distributed to expert , while denotes the proportion of the gating probability assigned to expert . To ensure an even distribution of the batch of tokens across the experts, the load-balancing loss function should be minimized. The optimal condition, i.e., , is achieved when each expert receives an equal number of dispatched tokens , and an equal proportion of the gating probability . The balance is thus maintained across all experts, ensuring that the workload is uniformly distributed at all times. Throughout the subsequent sections, unless explicitly stated otherwise, the term “MoE” will refer to “sparse MoE”.
3. Taxonomy of Mixture of Experts
@@
To effectively scale model parameters without a corresponding increase in computational demand, the mixture of experts (MoE) architecture has emerged as a viable solution. MoE leverages a collection of specialized models and a gating mechanism to dynamically select the appropriate “expert networks” for processing a given input. This enables the model to allocate computational resources on an as-needed basis, a concept known as conditional computation. The incorporation of MoE architectures into large language models (LLMs) is now a prevalent practice, allowing these models to achieve significant parameter scale-ups and consequent enhancements in capabilities (Fedus et al., 2022a; Lepikhin et al., 2020; Fedus et al., 2022b; Jiang et al., 2024a; Tianwen Wei, 2024).
forked edges,
for tree=
grow=east,
reversed=true,anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
draw=line-color,
rounded corners,align=left,
minimum width=2.5em,
s sep=3pt,
inner xsep=2pt,
inner ysep=1pt,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=3.2em,font=,,
where level=2text width=3.85em,font=,
where level=3text width=3.48em,font=,where level=4text width=3.8em,font=,[Mixture of Experts (MoE), ver
[Algorithm
[Gating
Function
[Dense
[DS-MoE(Pan et al., 2024), EvoMoE(Nie et al., 2021), MoLE(Wu et al., 2023b), LoRAMoE(Dou et al., 2023b)
,leaf,text width=20.2em]
[Sparse
[Token-Choice
[Shazeer et al.(Shazeer et al., 2017), GShard(Lepikhin et al., 2020), Switch Transformer(Fedus et al., 2022b)
ST-MoE(Zoph et al., 2022), M6-t(Yang et al., 2021), Mod-Squad(Chen et al., 2023a), StableMoE(Dai et al., 2022)
ModuleFormer(Shen et al., 2023c), OpenMoE(Xue et al., 2024), Skywork-MoE (Tianwen Wei, 2024)
Base Layer(Lewis et al., 2021), DSelect-k(Hazimeh et al., 2021), V-MoE(Riquelme et al., 2021), Z-code M3(Kim et al., 2021)
S-Base(Clark et al., 2022), Sentence-level MoE(Kudugunta et al., 2021), NLLB(Costa-jussà et al., 2022), MMoE(Ma et al., 2018)
Task-level MoE(Ye et al., 2022), X-MoE(Chi et al., 2022), Uni-Perceiver-MoE(Zhu et al., 2022)
Mixtral-8x7B(Jiang et al., 2024a), DeepSeekMoE(Dai et al., 2024), Jamba(Lieber et al., 2024), DBRX(Databricks, 2024),
MoA(Zhang et al., 2022b), JetMoE (Shen et al., 2024), Yuan 2.0-M32 (Wu et al., 2024b), DS-MoE(Pan et al., 2024)
,leaf,text width=15em]
[Non-trainable
Token-Choice
[Hash Layer(Roller et al., 2021), THOR(Zuo et al., 2021), DEMix(Gururangan et al., 2022), Task-MoE(Kudugunta et al., 2021)
M2M-100(Fan et al., 2021), Pangu-(Ren et al., 2023)
,leaf,text width=15em]
[Expert-Choice
[Expert-Choice MoE(Zhou et al., 2022a), Brainformers(Zhou et al., 2023)
,leaf,text width=15em]
[Soft
[Token Merging
[Soft MoE(Puigcerver et al., 2023), HOMOE(Dat et al., 2023)
,leaf,text width=15em]
[Expert Merging
[SMEAR(Muqeeth et al., 2023), Lory(Zhong et al., 2024), MoV(Zadouri et al., 2023), Omni-SMoLA(Wu et al., 2023a)
,leaf,text width=15em]
[Experts
[Network
Types
[FFN
[GShard(Lepikhin et al., 2020), Switch Transformer(Fedus et al., 2022b), ST-MoE(Zoph et al., 2022)
Branch-Train-MiX(Sukhbaatar et al., 2024), DS-MoE(Pan et al., 2024), MoEfication(Zhang et al., 2022a)
,leaf,text width=15em]
[Attention
[MoA(Zhang et al., 2022b), JetMoE(Shen et al., 2024), DS-MoE(Pan et al., 2024), ModuleFormer(Shen et al., 2023c)
,leaf,text width=15em]
[Others
[pMoE(Chowdhury et al., 2023), ADVMOE(Zhang et al., 2023), Chen et al.(Chen et al., 2022a), DeepMoE(Wang et al., 2020)
,leaf,text width=15em]
[Hyper-
parameters
[Count
[GShard(Lepikhin et al., 2020), Meta-MoE(Artetxe et al., 2021), GLaM(Du et al., 2022), Mixtral-8x7B(Jiang et al., 2024a)
ST-MoE(Zoph et al., 2022), Swith Transformer(Fedus et al., 2022b), DeepSpeed-MoE(Rajbhandari et al., 2022)
,leaf,text width=15em]
[Size
[GLaM(Du et al., 2022), DeepSeekMoE(Dai et al., 2024), DeepSeek-V2(DeepSeek-AI, 2024), DBRX(Databricks, 2024)
Qwen1.5-MoE(Team, 2024b), LLAMA-MoE (Team, 2023), Mixtral-8x7B(Jiang et al., 2024a)
,leaf,text width=15em]
[Frequency
[ST-MoE(Zoph et al., 2022), OpenMoE(Xue et al., 2024), V-MoE(Riquelme et al., 2021), MoE-LLaVA(Lin et al., 2024)
DeepSeekMoE(Dai et al., 2024), Brainformers(Zhou et al., 2023), Jamba(Lieber et al., 2024)
,leaf,text width=15em]
[Activ. Func.
[ReLU(Glorot et al., 2011), GeLU(Hendrycks and Gimpel, 2016), GeGLU(Shazeer, 2020), SwiGLU(Shazeer, 2020)
,leaf,text width=20.2em]
[Share Expert
[DeepSpeed-MoE(Rajbhandari et al., 2022), NLLB(Costa-jussà et al., 2022), DeepSeekMoE(Dai et al., 2024), OpenMoE(Xue et al., 2024), ScMoE(Cai et al., 2024)
Qwen1.5-MoE(Team, 2024b), MoCLE(Gou et al., 2023), Snowflake-Arctic(Team, 2024c)
,leaf,text width=20.2em]
[PEFT
[FFN
[LoRAMoE(Dou et al., 2023b),
AdaMix(Wang et al., 2022),
MixDA(Diao et al., 2023),
LLaVA-MoLE(Chen et al., 2024)
MixLoRA(Li et al., 2024b)
,leaf,text width=15em]
[Attention
[MoELoRA(Luo et al., 2024),
MoCLE(Gou et al., 2023),
SiRA(Zhu et al., 2023b)
,leaf,text width=15em]
[Transformer
Block
[MoV(Zadouri et al., 2023),
MoLoRA(Zadouri et al., 2023),
UniPELT(Mao et al., 2022),
Omni-SMoLA(Wu et al., 2023a)
MoLA(Gao et al., 2024),
MOELoRA(Liu et al., 2023)
,leaf,text width=15em]
[Every Layer
[MoLE(Wu et al., 2024a)
,leaf,text width=15em]
[Training
& Inference
Scheme
[Original
[Shazeer et al.(Shazeer et al., 2017), GShard(Lepikhin et al., 2020), Switch Transformer(Fedus et al., 2022b), ST-MoE(Zoph et al., 2022)
,leaf,text width=20.2em]
[Dense2Sparse
[RMoE(Wu et al., 2022), Dua et al.(Dua et al., 2022), Sparse Upcycling(Komatsuzaki et al., 2022), DS-MoE(Pan et al., 2024), EvoMoE(Nie et al., 2021)
MoE-LLaVA(Lin et al., 2024), SMoE-Dropout(Chen et al., 2022d), Skywork-MoE(Tianwen Wei, 2024), MoEfication (Zhang et al., 2022a)
MoEBERT(Zuo et al., 2022), LLaMA-MoE(Team, 2023)
,leaf,text width=20.2em]
[Sparse2Dense
[OneS(Xue et al., 2022a), MoE-Pruning(Chen et al., 2022b), ModuleFormer (Shen et al., 2023c), EWA(Huang et al., 2023)
,leaf,text width=20.2em]
[Expert Models
Merging
[Branch-Train-Merge(Li et al., 2022), Branch-Train-MiX(Sukhbaatar et al., 2024), FoE(Wang et al., 2023)
,leaf,text width=20.2em]
[Derivatives
[Lifelong-MoE(Chen et al., 2023c), MoT(Antoniak et al., 2023), MoD(Raposo et al., 2024), WideNet(Xue et al., 2022b), SUT(Tan et al., 2023), SMoP(Choi et al., 2023)
,leaf,text width=25.2em]
[System
[Computation
[FastMoE(He et al., 2021), DeepSpeed-MoE(Rajbhandari et al., 2022), Tutel(Hwang et al., 2023a), SE-MoE(Shen et al., 2022), FasterMoE(He et al., 2022), DeepSpeed-TED(Singh et al., 2023)
HetuMoE(Nie et al., 2022), FlexMoE(Nie et al., 2023), SmartMoE(Zhai et al., 2023), MegaBlocks(Gale et al., 2023),
ScatterMoE(Tan et al., 2024),
PIT(Zheng et al., 2023)
,leaf,text width=25.2em]
[Communication
[DeepSpeed-MoE(Rajbhandari et al., 2022), HetuMoE(Nie et al., 2022), FasterMoE(He et al., 2022), ExFlow(Yao et al., 2024), Tutel(Hwang et al., 2023a), DeepSpeed-TED(Singh et al., 2023)
TA-MoE(Chen et al., 2022c), SE-MoE(Shen et al., 2022),
MPipeMoE (Zhang et al., 2024), SkyWork-MoE(Tianwen Wei, 2024), Lancet(Jiang et al., 2024b), ScMoE(Cai et al., 2024), Arctic(Team, 2024c)
,leaf,text width=25.2em]
[Storage
[SE-MoE(Shen et al., 2022), Pre-gated MoE(Hwang et al., 2023b), EdgeMoE(Yi et al., 2023), MPipeMoE(Zhang et al., 2024)
,leaf,text width=25.2em]
[Application
[NLP
[Shazeer et al.(Shazeer et al., 2017), GShard(Lepikhin et al., 2020), Swith Transformer(Fedus et al., 2022b), Meta-MoE(Artetxe et al., 2021), GLaM(Du et al., 2022), NLLB(Costa-jussà et al., 2022)
Mixtral-8x7B(Jiang et al., 2024a) DeepSpeed-MoE(Rajbhandari et al., 2022), ST-MoE(Zoph et al., 2022), DeepSeekMoE(Dai et al., 2024)
,leaf,text width=25.2em]
[CV
[V-MoE(Riquelme et al., 2021), Swin-MoE(Hwang et al., 2023a), ADVMOE(Zhang et al., 2023), pMoE(Chowdhury et al., 2023)
,leaf,text width=25.2em]
[RecSys
[MMoE(Ma et al., 2018), PLE(Tang et al., 2020), AdaMCT(Jiang et al., 2023b)
,leaf,text width=25.2em]
[MultiModal
[LIMoE(Mustafa et al., 2022), Shen et al.(Shen et al., 2023b), MoCLE(Gou et al., 2023), LLaVA-MoLE(Chen et al., 2024), MoE-LLaVA(Lin et al., 2024), Uni-MoE(Li et al., 2024a)
MM1(McKinzie et al., 2024)
,leaf,text width=25.2em]
For example, the Mixtral 8x7B (Jiang et al., 2024a), introduced by Mixtral AI, shares its foundational architecture with the earlier Mistral 7B (Jiang et al., 2023a), but with a notable difference: each layer comprises eight feed-forward networks (FFN) (i.e., experts). Despite utilizing only 13 billion active parameters, the Mixtral-8x7B demonstrates superior or equivalent performance to the Llama-2-70B (Touvron et al., 2023) and GPT-3.5 (OpenAI, 2022) across various benchmarks. Similarly, the DeepSeek LLM (Bi et al., 2024), developed by DeepSeek, has been extended with an MoE variant known as DeepSeekMoE (Dai et al., 2024). The DeepSeekMoE 16B, while requiring approximately 40% less computation, attains performance on par with the Llama 2 7B (Touvron et al., 2023). The Qwen team has also contributed to this innovative field by developing the Qwen1.5-MoE (Team, 2024b), a smaller MoE model with only 2.7B active parameters that rivals the performance of leading 7B parameter models such as the Mistral 7B (Jiang et al., 2023a) and the Qwen1.5-7B (Team, 2024a).
To assist researchers in navigating the rapidly evolving landscape of LLMs equipped with MoE architectures, we have developed a taxonomy that categorizes these models from three perspectives: algorithm design, system design, and application. Figure 3 showcases our taxonomy alongside several representative studies. In the following sections, we will provide a comprehensive and in-depth analysis of each category within our taxonomy.
4. Algorithm Design of Mixture of Experts
4.1. Gating Function

### v2

@@ Abstract
Large language models (LLMs) have garnered unprecedented advancements across diverse fields, ranging from natural language processing to computer vision and beyond. The prowess of LLMs is underpinned by their substantial model size, extensive and diverse datasets, and the vast computational power harnessed during training, all of which contribute to the emergent abilities of LLMs (e.g., in-context learning) that are not present in small models. Within this context, the mixture of experts (MoE) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead, gaining significant attention from academia and industry. Despite its growing prevalence, there lacks a systematic and comprehensive review of the literature on MoE. This survey seeks to bridge that gap, serving as an essential resource for researchers delving into the intricacies of MoE. We first briefly introduce the structure of the MoE layer, followed by proposing a new taxonomy of MoE. Next, we overview the core designs for various MoE models including both algorithmic and systemic aspects, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, we delineate the multifaceted applications of MoE in practice, and outline some potential directions for future research. To facilitate ongoing updates and the sharing of cutting-edge advances in MoE research, we have established a resource repository accessible at https://github.com/withinmiaov/A-Survey-on-Mixture-of-Experts.
Index Terms:
Large Language Models, Mixture of Experts, Gating Functions1 Introduction
@@
In the current landscape of artificial general intelligence (AGI), the transformative impact of transformer-based large language models (LLMs) has permeated diverse fields such as natural language processing [1, 2, 3, 4, 5, 6, 7], computer vision [8, 9], and multimodality [10, 11, 12, 13]. Building upon the foundational transformer architecture, LLMs demonstrate extraordinary capabilities, which are attributed to their sheer size, the breadth of data they are trained on, and the significant computational resources invested in their development [14, 15, 16]. Recognizing a scaling law [14, 17] that underpins their evolution, it is imperative to identify and implement efficient methodologies for the sustainable scaling of LLMs.
@@
The concept of mixture of experts (MoE), initially introduced in [18, 19], has undergone extensive exploration and advancement as evidenced by subsequent studies [20, 21, 22, 23, 24, 25, 26]. The emergence of sparsely-gated MoE [27], particularly within the integration of transformer-based large language models [28], has brought new vitality to this three-decade-old technology. The MoE framework is based on a simple yet powerful idea: different parts of a model, known as experts, specialize in different tasks or aspects of the data. With this paradigm, only pertinent experts are engaged for a given input, keeping the computational cost in check while still benefiting from a large pool of specialized knowledge. This scalable and flexible innovation has offered an effective approach for adhering to the scaling law, allowing for increased model capacity without a corresponding surge in computational demands. As depicted in Figure 1, MoE has maintained a robust trajectory of growth, particularly notable in 2024 with the advent of Mixtral-8x7B [29] and a variety of subsequent industrial-scale LLMs such as Grok-1 [30], DBRX [31], Arctic [32], DeepSeek-V2 [33], etc.
@@
Despite the increasing popularity and application of MoE models in various domains, the literature has yet to see a survey that thoroughly examines and categorizes the advancements in this area. The most recent review of MoE we could find was presented in September 2022 [34], predating the pivotal “ChatGPT moment”, which omits the significant advancements that have recently emerged alongside the escalating academic and industrial interest in this domain. This gap in the literature not only hinders the progress of MoE research but also limits the dissemination of knowledge on this topic to a broader audience. Our survey aims to address this deficit by providing a clear and comprehensive overview of MoE with a novel taxonomy that segments recent progress into algorithm, system, and application.
@@
Under this taxonomy, we first delve into MoE algorithmic advancements, particularly the prevalent substitution of feed-forward network (FFN) layers with MoE layers in transformer-based LLMs [28, 35, 36, 37, 29, 38, 33]. As each MoE layer integrates multiple FFNs—each designated as an expert—and employs a gating function to activate a selected subset of these experts, we explore the design choices of gating function and expert network, alongside collections of available open-source implementations, hyperparameter configurations, and empirical evaluations. Furthermore, to underscore the flexibility and versatility of MoE, we extend our analysis beyond the standard integration of MoE into model backbone, and discuss an array of novel MoE-related designs, such as soft MoE with token or expert merging [39, 40, 41, 42, 43], mixture of parameter-efficient experts (MoPEs) [44, 42, 45, 46, 47, 48], training and inference schemes with model transition between dense and sparse [49, 50, 51, 52, 53, 54], and various derivatives [55, 56, 57, 58, 59, 60].
@@
With the gradual convergence of model architecture design in industrial products, system design has emerged as a pivotal factor in enhancing the quality of LLM services. Given the close association of MoE models with machine learning system design, we provide a comprehensive overview of MoE system design, including computation, communication, and storage enhancements tailored to address the unique challenges posed by the sparse and dynamic nature of its computational workload. Additionally, we overview the applications of MoE across various domains, including natural language processing, computer vision, recommender system, and multimodal contexts.
@@
The remainder of this survey is organized as follows. Section 2 provides a foundational understanding of MoE, contrasting sparse and dense activation of experts. Section 3 introduces our proposed taxonomy for categorizing MoE advancements. Sections 4, 5, and 6 delve into the algorithmic designs, computing system support, and various applications of MoE models, respectively, following the structure outlined in our taxonomy in Figure 3. Finally, in Section 7, we highlight the critical challenges and opportunities for bridging the research-practicality gap, culminating in Section 8 with our conclusions.
2 Background on Mixture of Experts
@@
In transformer-based large language models (LLMs), each mixture-of-experts (MoE) layer typically consists of a set of “expert networks” , alongside a “gating network” . The role of the gating network, which often takes the form of a linear network with a softmax activation function, is to direct the input to the appropriate expert networks [27, 36]. The MoE layer is strategically placed to select the feed-forward network (FFN) within each Transformer block, typically following the self-attention (SA) sub-layer. This positioning is crucial because the FFN become increasingly computationally demanding as the model scales up. For instance, in the PaLM [5] model with the parameter number of 540B, the 90% of these parameters are within its FFN layers.
@@
Formally, each expert network , usually a linear-ReLU-linear network, is parameterized by , accepting the same input and generating an output . In parallel, the gating network , parameterized by and typically consisting of a linear-softmax network, yields the output . Based on the design of gating function, MoE layers can be broadly classified into two categories: dense MoE and sparse MoE, which are described in detail in the following subsections.
2.1 Dense MoE
@@
The dense mixture-of-experts layer activates all expert networks during each iteration. This strategy has been extensively employed in a range of early proposals [18, 19, 20, 21, 22, 23, 24, 25, 26, 61]. Most recently, the dense MoE concept has been revisited by studies such as EvoMoE[62], MoLE [63], LoRAMoE [45], and DS-MoE[64]. The structure of the dense MoE layer is depicted in Figure 2(a). Consequently, the output of the dense MoE layer can be formulated as
@@ 2.2 Sparse MoE
While a dense mixture of experts typically yields higher prediction accuracy [8], it also incurs a significant increase in computational overhead. To address this, Shazeer et al. [27] introduced the sparsely-gated MoE layer, which is designed to activate only a selected subset of experts during each forward pass. This strategy achieves sparsity by computing a weighted sum of the outputs from only the top- experts, rather than aggregating the outputs from all the experts. The structure of the sparse MoE layer is illustrated in Figure 2(b). Building on the framework established by [27], Equation (2) can be modified to reflect the sparsely-gated mechanism as follows:
@@
To explain, function retains only the top- entries of a vector at their original values, while setting all other entries to . Following the operation, those entries assigned become approximately zero. The hyper-parameter is selected based on the specific application, with common choices being [36, 65] or [28, 35, 66, 37, 29, 67]. The addition of a noise term is a prevalent strategy for training a sparsely-gated MoE layer, fostering exploration among experts and enhancing the stability of MoE training [27, 36].
@@
Although the sparse gate substantially expands the model’s parameter space without a corresponding increase in computational cost, it can lead to a load balancing issue. Such an issue refers to the uneven distribution of workload across experts, with some being frequently utilized and others seldom or never engaged. To address this, each MoE layer incorporates an auxiliary loss function that promotes an even distribution of tokens across experts within each batch, as described in many studies [28, 36, 35, 29, 68, 69, 67]. To formulate this concept, consider a batch of queries , comprising tokens, and experts indexed from to . Following [28, 36], the auxiliary load balancing loss for the batch is defined as
@@
where represents the proportion of tokens distributed to expert , while denotes the proportion of the gating probability assigned to expert . To ensure an even distribution of the batch of tokens across the experts, the load-balancing loss function should be minimized. The optimal condition, i.e., , is achieved when each expert receives an equal number of dispatched tokens , and an equal proportion of the gating probability . The balance is thus maintained across all the experts, ensuring that the workload is uniformly distributed at all times. Throughout the subsequent sections, unless explicitly stated otherwise, the term “MoE” refers to “sparse MoE”.
3 Taxonomy of Mixture of Experts
@@
To effectively scale model parameters without a corresponding increase in computational demand, the mixture of experts (MoE) architecture has emerged as a viable solution. MoE leverages a collection of specialized models and a gating mechanism to dynamically select the appropriate “expert networks” for processing a given input. This enables the model to allocate computational resources on an as-needed basis, a concept known as conditional computation. The incorporation of MoE architectures into large language models (LLMs) is now a prevalent practice, allowing these models to achieve significant parameter scale-ups and consequent enhancements in capabilities [34, 28, 36, 29, 67].
For example, the Mixtral 8x7B [29], introduced by Mixtral AI, shares its foundational architecture with the earlier Mistral 7B [160], but with a notable difference: each layer comprises eight feed-forward networks (FFN) (i.e., experts). Despite utilizing only 13 billion active parameters, the Mixtral-8x7B demonstrates superior or equivalent performance to the Llama-2-70B [161] and GPT-3.5 [162] across various benchmarks. Similarly, the DeepSeek LLM [163], developed by DeepSeek, has been extended with an MoE variant known as DeepSeekMoE [69]. The DeepSeekMoE 16B, while requiring approximately 40% less computation, attains performance on par with the Llama 2 7B [161]. The Qwen team has also contributed to this innovative field by developing the Qwen1.5-MoE [104], a smaller MoE model with only 2.7B active parameters that rivals the performance of leading 7B parameter models such as the Mistral 7B [160] and the Qwen1.5-7B [164].
To assist researchers in navigating the rapidly evolving landscape of LLMs equipped with MoE architectures, we have developed a taxonomy that categorizes these models from three perspectives: algorithm design, system design, and application. Figure 3 showcases our taxonomy alongside several representative studies. In the following sections, we provide a comprehensive and in-depth analysis of each category within our taxonomy.
4 Algorithm Design of Mixture of Experts
4.1 Gating Function